### v2

# A fast algorithm for 2D Rigidity Percolation

Abstract
Rigidity Percolation is a crucial framework for describing rigidity transitions in amorphous systems. We present a new, efficient algorithm to study central-force Rigidity Percolation in two dimensions. This algorithm combines the Pebble Game algorithm, the Newman-Ziff approach to Connectivity Percolation, as well as novel rigorous results in rigidity theory, to exactly identify rigid clusters over the full bond concentration range, in a time that scales as $N^{1.02}$ for a system of $N$ nodes. We perform extensive numerical simulations with systems larger than $500$ million nodes, far beyond the previous limitations. We obtain new, precise estimates for the critical exponents, $\nu=1.1694(8)$ and $D_{f}=1.8423(7)$, and locate the critical threshold at $p_{c}=0.6602741(4)$. Besides opening the way to further accurate numerical studies of Rigidity Percolation, our work provides new rigorous theoretical insights on specific cluster merging mechanisms that distinguish it from the standard Connectivity Percolation problem.
preprint: APS/123-QED
I Introduction
Percolation is one of the most versatile conceptual frameworks in statistical physics and describes a wealth of natural phenomena, from polymer gelation to the propagation of diseases. In standard –Connectivity– Percolation (CP) the central objects are connectivity clusters, namely sets of mutually connected nodes. As more and more bonds are added to the network, clusters grow and coalesce, and a system-spanning (macroscopic) cluster eventually emerges at a critical value of the bond concentration.
While CP theory provides an extensive framework to characterize the connectivity properties of networks [1, 2], Rigidity Percolation (RP) addresses instead their mechanical stability: is the network able to transmit stresses and sustain external loads? It was realized early that this question can be formulated as a constraint-counting problem [3, 4]; the rigidity of a network depends only on its topology and is largely independent of the precise physical nature of bonds. In central-force networks, on which we focus in this work, nodes possess $d$ degrees of freedom in $d$ dimensions, while bonds introduce constraints by fixing the distances between pairs of nodes. Rigid clusters are sets of mutually rigid bonds, that is, subnetworks in which all degrees of freedom are fixed by the constraints except for the $d(d+1)/2$ global ones. Floppy clusters instead, have unfixed degrees of freedom that allow for deformation along soft (floppy) modes. While at low bond concentration the system is typically macroscopically floppy (i.e. composed of small, non-rigidly connected rigid clusters), increasing the concentration above a critical threshold leads to the emergence of a macroscopic rigid cluster (see e.g. [5]). Rigidity Percolation has attracted much attention in Soft Matter, as a simple and versatile framework to model liquid to amorphous solid transitions in systems such as colloidal gels [6, 7, 8], living tissues [9, 10, 11, 12], fiber networks [13, 14] or granular packings [15, 16, 17, 18]. Notably, it has recently been revealed how the RP critical exponents control the viscoelastic properties of colloidal gels in their solid phase [19, 20]. Moreover, recent works gather evidence for a larger role of rigid clusters in the physics of amorphous systems, as controlling e.g. the behaviour at the brittle-ductile transition [21] or the large stress fluctuations at the shear thickening transition [22].
Inquiring about mutual rigidity, rather than mutual connectivity, makes RP much more difficult than the usual CP problem. A distinguishing feature of RP is its non-locality: as we make precise in section IV, the activation of a single bond can trigger the rigidification of arbitrarily large spatial regions, by coalescence of numerous rigid clusters into a single one. Likewise, cutting a single bond can lead to cascades of rigidity loss [23]. As a matter of fact, RP remains largely less understood than CP. Rigorous results are rather scarce: the existence of a critical bond concentration strictly above the CP threshold has been proven on the 2D triangular lattice [24]; the uniqueness of the infinite rigid cluster has been proven in Ref. [25]. Exact results have been obtained on non-generic networks such as hierarchical and Erdös-Rényi graphs [26, 27] (see also [28, 29] for connections with satisfiability problems). In most cases however, especially those relevant to Soft Matter applications, the characterization of the RP universality classes relies essentially on numerical simulations, either to determine the values of critical exponents [5, 16], or to explore fine aspects of the universality class such as conformal invariance [30, 31].
Yet, identifying efficiently rigid clusters in large systems is not an easy task. Numerical studies were made possible thanks to the Pebble Game algorithm [32, 5] (later generalized to frictional systems [15, 33]), a combinatorial implementation of Laman theorem for graph rigidity, that we briefly describe in section II. Using the Pebble Game algorithm, the time to identify rigid clusters at $p_{c}^{RP}$ scales with the system size $N$ as $N^{1.15}$ [32]. This allowed to estimate the values of the critical exponents for central-force RP, $\nu^{\rm RP}=1.21\pm 0.06,\;\beta^{\rm RP}=0.175\pm 0.02,\;D_{f}^{\rm RP}=1.86\pm 0.02$, respectively the correlation length and order parameter exponents, and the fractal dimension of the spanning rigid cluster, with a sufficient precision to claim the existence of a new universality class, distinct from random CP [5]. To our knowledge, there has been no attempt to determine the central-force RP exponents with better precision since then, the barrier being the harsh computational cost of analyzing systems larger than $\sim 10^{6}$ nodes (for comparison, accurate studies of the CP critical point involve systems with $O(10^{8})$ nodes [34]).
Classifying rigidity transitions as belonging or not to the (uncorrelated  [^1] In our, as well as earlier work [5], bonds are activated in a random, uncorrelated way. Recent work has examined the effect of local correlations [6] but to our knowledge the effect of long-range correlations in the bond activation on the RP universality class has not been addressed. Our algorithm can be generalized to incorporate such correlations and we plan this for future work.) central-force RP universality class is crucial to achieve a better understanding of the variety of liquid-to-amorphous solid transitions [6, 36, 9, 37, 16, 38, 39], and to assess the robustness of the central-force RP universality class. Providing accurate numerical standards for the critical exponents, and, more generally, understanding the specificities of the RP problem as compared to the well-studied CP one, are therefore major goals.
In this paper we introduce a new algorithm for RP. It is based on the powerful Newman-Ziff (NZ) approach to CP[40], where bonds are activated one at a time, and the state of the system, namely the identification of clusters, is fully known at each step. Studying percolation à la Newman-Ziff allows notably to measure the quantities of interest (probability to percolate, size of the largest component) at each bond activation, namely for each value of the bond filling fraction (fraction of active bonds), thereby covering the complete phase diagram in a single run. To obtain a NZ-like algorithm for RP, we first shed new light on the specific cluster merging mechanisms at play, by providing rigorous results that predict the consequences of each new bond activation. Exploiting these results we construct an algorithm that scales with the system size as $N^{1.02}$, i.e. that covers the entire phase diagram in almost linear time. Our algorithm pushes the size limitation to $N\gtrsim 10^{8}$, allowing for precise studies of the RP universality class, starting with accurate determination of the main critical exponents [41].
The paper is organized as follows. In section II we briefly describe the standard Pebble Game as defined by Jacobs and Hendrickson [32, 5], and explain in more details how to perform pebble searches, that are key to the implementation of our own algorithm. In section III we recall the Newman-Ziff algorithm for CP. In section IV we present the three theorems that allow us to design a NZ-like algorithm for RP, namely that allow, with minimal computational cost, to deduce the new state of the system after each bond activation. The proofs of those theorems, as well as a brief review of 2D rigidity theory and additional new results, are presented in the Supplemental Material (SM), section 1. We give an overview of the strategy behind our algorithm in section V, while the implementation is detailed in section VI. In section VII we adapt the (CP) Machta algorithm [42] to detect the wrapping of rigid clusters. Finally we discuss the performance of our complete algorithm in section VIII. We conclude with several perspectives in section IX.
II The Jacobs and Hendrickson approach
The central challenge in studying RP is the identification of rigid clusters, namely of the maximal sets of mutually rigid bonds. In 2D, an exact criterion is given by Laman theorem on graph rigidity [4] (cf. SM section 1 for more details). Jacobs and Hendrickson (JH) [32] introduced a clever algorithm to implement efficiently Laman’s criterion, the Pebble Game: the degrees of freedom of the system are represented as “pebbles”, assigned to the nodes. Pebbles are removed each time an independent bond (constraint) is added to the system. Note that, in this formulation, three pebbles can always be gathered across any bond, since the system always has three global degrees of freedom in 2D. Additional pebbles correspond to floppy modes.
II.1 The Pebble Game
The JH algorithm is divided into two stages, filling and clustering. All bonds are initially inactive and each node has two pebbles attached to it, representing its two degrees of freedom. In the filling stage, bonds are independently activated with probability $p$. Each new activated bond is classified as an independent or redundant constraint using pebble searches (see next section): a new bond $uv$ represents an independent constraint if $u$ and $v$ were not mutually rigid, namely if four pebbles can be gathered across $uv$. If the two nodes were already mutually rigid, only three pebbles can be gathered, and the new bond is redundant.
Once all the bonds are activated and classified, rigid clusters are identified. Each rigid cluster is constructed by starting from a given unlabeled bond $b$, and testing the mutual rigidity of the other (unlabeled) bonds of the system with $b$, by performing pebble searches. Each bond that is found to be mutually rigid with respect to $b$, is given the same label as $b$. This is conveniently implemented as a Breadth First Search that stops when only floppy, or already labeled bonds, can be found.
By construction, one run of the JH algorithm corresponds to a single value of the bond concentration $p$. The overall cost of the JH algorithm is determined by the number of pebble searches to be performed in a run –which depends on NN– as well as the cost of each search –which also depends on $N$. JH reported a scaling of the algorithm that depends on $p$, scaling as $N^{1.15}$ near the critical point $p\approx p_{c}$, and linearly near full concentration. In practice this means that studying the phase diagram with the JH algorithm requires to run it for (as many as possible) different values of $p$, with a total computing time that depends on the number of values considered, most of which being typically concentrated near the critical point with cost $N^{1.15}$ each.
Our algorithm solves these two limitations at once. First, the whole phase diagram is obtained in a single run. This comes from adopting a Newman-Ziff approach, which we describe in the next section. Secondly, we make such approach efficient by exploiting rigidity theory to make a strategic use of pebbles searches, keeping their number low.
II.2 Pebble searches
Pebbles searches are used to check for mutual rigidity. These searches take place on the pebble graph, distinct from the physical lattice. Everytime an independent bond of the system is activated, a directed edge is added to the pebble graph and a pebble is removed from one of the end nodes, accounting from the removal of one degree of freedom. The edge is initially directed away from the node that loses a pebble. For later use, we distinguish two types of pebble searches, whose dynamics is illustrated in Fig. 1.
Type I searches – The purpose of pebble searches of type I is to actually move the pebbles across the pebble graph, in order to keep track of the repartition of degrees of freedom. To gather a pebble at node $j$, a Breadth First Search (BFS)  [^2] Depth First Search can be used as well. JH used BFS and we do the same in our work. starting at $j$ is performed over the pebble graph, with the aim of finding a path that leads from node $j$ to a node that has a pebble (Fig. 1a). If a pebble is found, the direction of the pebble edges along the path is reversed and the pebble is moved from the node where it has been found to node $j$ (Fig. 1b).
Type II searches – The purpose of pebble searches of type II is to test mutual rigidity without actually moving the pebbles across the pebble graph. To check mutual rigidity of bonds $b_{1}$ and $b_{2}$, three pebbles (that are always available) are first gathered at $b_{1}$ using type I searches. These pebbles are temporarily frozen at their location, i.e., they are made unavailable to the rest of the system, while the nodes of $b_{1}$ are marked as rigid. Then, a search is performed for a fourth pebble that could be gathered at the nodes of $b_{2}$: if such pebble exists, bonds $b_{1}$ and $b_{2}$ are mutually floppy, otherwise they are mutually rigid. Note that, for $b_{2}$ to be rigid with respect to $b_{1}$, both its nodes must be rigid. If one node is rigid and the other is floppy, the bond is overall floppy. This implies that, in contrast to CP, nodes in RP can belong to more than a rigid cluster. Each bond, however, belongs to a single rigid cluster. Nodes belonging to more than a rigid cluster are called pivots [32].
The search proceeds as in the previous section, but the direction of pebble edges is never changed by type II searches, as pebbles are not actually moved. Moreover, if the search is successful (a pebble turns out to be available), all nodes along the path are marked as floppy (Fig. 1e). Indeed, the found pebble could be gathered at any of them. On the contrary, after a failed search, all the visited nodes are marked rigid (Fig. 1f): no pebble can be gathered at any of those visited nodes.
This marking of the nodes reduces the cost of subsequent searches, that can be terminated whenever a node, already marked as floppy during a previous search, is encountered. Moreover, edges of the pebble graph that start at nodes marked rigid during previous searches are not walked, as they cannot lead to any pebble nor to a floppy node [32].
Triangulation – Note that both types of searches can, in principle, result in a pebble not being found. When a pebble search fails, all the pebble edges traversed during the search are triangulated over the bond that has three pebbles (Fig. 1c). The triangulation procedure is essential to keep pebble searches short [32].
Figure 1: Dynamics of pebble searches on the pebble graph. Top: searches of type I. Bottom: searches of type II. To ease visualization, bonds are not triangulated in the bottom panels. (a) Green nodes have two pebbles, blue nodes have one, grey have zero. Pebbles are depicted as black dots. The red undirected bond is the newly activated bond. A pebble search of type I starts at the grey end node of the new bond. The purple path leads to a pebble. (b) The path identified in (a) is reversed, the node where it started has now one pebble, and a second search is started. The search traverses the black edges without finding a pebble. (c) The edges traversed during the previous search are triangulated over the base made by the new bond. (d) Red (blue) nodes have been marked rigid (floppy) during previous searches, while open circles represent unmarked nodes. A pebble search of type II starts at the yellow star. The search traverses the black edges without finding a floppy node. Note that edges originating from rigid nodes are not walked. (e) The nodes visited during the search of panel (d) are marked rigid. The visited edges are triangulated (not shown). A new search starts at the yellow star and, following the blue path, stops when it hits a floppy node. (f) The nodes traversed by the path identified in (e) are marked floppy.
III The Newman-Ziff algorithm for CP
In this section we recall briefly the NZ approach to CP [44, 40]. NZ simulated percolation as a dynamical process, where bonds are activated one by one, and the state of the system is updated at each step, following simple rules. The algorithm allows therefore to generate instances of the system at every filling fraction, and does so in linear time. Measurements are then convoluted with the binomial distribution to be transformed in functions of the bond concentration $p$ [40]. The key idea in the NZ algorithm for CP, which we adapt to RP, is to describe each cluster as a directed tree: all nodes belonging to a given connectivity cluster point to the same root in the corresponding tree. Each cluster is identified by its root, whose entry in the tree gives (minus) the size of the cluster.
When a new bond $uv$ is activated, trees (clusters) must be updated accordingly. These updates are simple in CP, as only two events can happen. If $u$ and $v$ already belong to the same cluster (they point to the same root), the corresponding tree is unchanged. If $u$ and $v$ belong to two distinct clusters, $\mathcal{C}_{u}$ and $\mathcal{C}_{v}$ (they point to different roots), the new bond coalesces them into $\mathcal{C}_{u}\cup\mathcal{C}_{v}\cup\,\{uv\}$. Coalescence is performed by redirecting the root of the smallest cluster toward the root of the largest one.
The roots of $u$ and $v$ are identified by the so-called find operation, i.e., by traversing the tree from $u$ ($v$) to the corresponding root. The find operation is coupled with the so-called path compression: once the root is found, the tree is compressed by redirecting all the traversed edges toward the identified root. This mechanism allows the find operation to have constant cost even on large trees [40].
IV Theorems for single bond activation
In this paper we adapt the NZ strategy described in the previous section, to rigidity percolation. Namely, starting with an empty lattice and progressively activating the bonds, we are able to deduce the new state of the system after each new bond activation. While this is relatively simple in CP, due to simple cluster events (cf. section III), the situation is more involved in RP, where the consequence of single bond activation is a priori not clear: is the new bond independent or redundant? Are new rigid clusters created or existing ones coalesced? If so, which ones and how? In this section we provide theoretical results in 2D rigidity theory that allow us to identify the only three possible cluster events resulting from single bond activation, that we name Pivoting, Overconstraining, and Rigidification. The following results allow us to furthermore determine, without making use of pebble searches, (i) if an added bond is independent or redundant, (ii) what type of cluster event is triggered by a given bond activation. In addition, defining the pivotal class of a node as the number of distinct rigid clusters the node belongs to  [^3] Inactive nodes have $\pi=0$ by definition., we also determine (iii) how the pivotal classes of nodes change. In section V we explain how to combine these results with optimal pebble searches to obtain our final algorithm. We start by giving basic principles of 2D rigidity, and refer to section 1 of the SM for further details.
Proposition 1 (Transitivity) – Denote $\mathcal{R}_{1}$, $\mathcal{R}_{2}$, $\mathcal{R}_{3}$ three rigid bodies.
  1. 1.
    If $\mathcal{R}_{1}$ is rigid with respect to $\mathcal{R}_{2}$ and $\mathcal{R}_{2}$ is rigid with respect to $\mathcal{R}_{3}$, then $\mathcal{R}_{1}$ is rigid with respect to $\mathcal{R}_{3}$.
  2. 2.
    If $\mathcal{R}_{1}$ is rigid with respect to $\mathcal{R}_{2}$ and $\mathcal{R}_{2}$ is floppy with respect to $\mathcal{R}_{3}$, then $\mathcal{R}_{1}$ is floppy with respect to $\mathcal{R}_{3}$.
From now on, $\mathcal{R}$ denotes a rigid cluster, $\mathcal{C}$ a connectivity cluster and $\mathcal{C}_{u}$ is the connectivity cluster that contains node $u$.
Proposition 2 (“Two-pivots rule”) [46] In two dimensions, if two rigid clusters $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ share at least two pivots, then $\mathcal{R}_{1}\cup\mathcal{R}_{2}$ is a rigid cluster.
In the following, $uv$ denotes a new activated bond. The proofs of the following statements are given in section 1B of the SM. Additional results on rigidity, which to our knowledge had not been proven before, are given in section 1A of the SM.
Theorem 1 (Pivoting) – Assume $u\in\mathcal{C}_{u}$ and $v\in\mathcal{C}_{v}$ with $\mathcal{C}_{u}\neq\mathcal{C}_{v}$. Then,
  1. 1.
    The bond $uv$ is independent.
  2. 2.
    The activation of $uv$ creates a new rigid cluster composed of the bond $uv$ only.
  3. 3.
    The pivotal class of $u$ and $v$ increases by 1. The pivotal class of any other node is unchanged.
Theorem 2 (Overconstraining) – Assume that there exists a rigid cluster $\mathcal{R}$ such that $u,v\in\mathcal{R}$ before the activation of $uv$. Then,
  1. 1.
    uvuv is a redundant bond.
  2. 2.
    Upon activation of $uv$, $\mathcal{R}$ becomes $\mathcal{R}\cup\{uv\}$; all other rigid clusters remain identical.
  3. 3.
    The pivotal classes of all nodes are unchanged.
Theorem 3 (Rigidification) – Assume $\mathcal{C}_{u}=\mathcal{C}_{v}=\mathcal{C}$ and that, before the activation of $uv$, there is no rigid cluster that contains both $u$ and $v$. Then,
  1. 1.
    uvuv is an independent bond.
  2. 2.
    The activation of $uv$ triggers a rigidification process resulting in the coalescence of $k+1$ rigid clusters $\mathcal{R}_{1},\,\mathcal{R}_{2},\cdots,\mathcal{R}_{k}$ and $\mathcal{R}_{k+1}=\{uv\}$ into a single rigid cluster $\mathcal{R}^{+}=\mathcal{R}_{1}\cup\mathcal{R}_{2}\cup\cdots\cup\mathcal{R}_{k}\cup\{uv\}$. The clusters $\mathcal{R}_{1},\,\mathcal{R}_{2},\cdots,\mathcal{R}_{k}$ all belong to $\mathcal{C}$, and each of them is connected to at least two other ones via at least two distinct pivots.
  3. 3.
    The pivotal class of these pivots decreases by at least one.
V Overview of the algorithm
In this section we present an algorithm that exploits the three theorems of section IV. How we efficiently implement it in practice is explained in the next section.
V.1 Basic settings
The system is initially empty, i.e., there are no active bonds or nodes: each node has pivotal class 0 and carries two pebbles. The pebble graph is empty.
For the hypotheses of the theorems to be verifiable, the connectivity state of the network must be known at each new bond activation. We therefore implement the NZ algorithm for CP as done in Ref. [40]. The NZ algorithm scales linearly with system size, so that, with a very small additional time and memory cost, our algorithm produces the entire phase diagram of the CP transition as well.
To monitor the rigidity state of the system we adopt the same strategy as NZ [44] and represent rigid clusters as trees. For the sake of clarity, we indicate the trees that describe connectivity and rigid clusters as $\mathcal{T}_{\rm{C}}$ and $\mathcal{T}_{\rm{R}}$ respectively. As nodes might belong to multiple rigid clusters, while bonds cannot, each node in a tree $\mathcal{T}_{\mathcal{R}}$ represents a bond in the corresponding rigid cluster $\mathcal{R}$. Traversing a tree $\mathcal{T}_{\rm{R}}$ leads to a root bond that uniquely identifies the rigid cluster. Every time a root is identified by a find operation, the path compression is also performed in order to keep a constant cost for each of these operations (cf. Section III).
V.2 Stages of the algorithm
As for the NZ algorithm for CP, our goal is to deduce the new state of the system after each new bond activation, namely, to update the rigid clusters trees from the knowledge of the system state prior to the bond activation. To this aim, we proceed as follows.
  1. 1.
    Select a random bond $uv$ to activate.
  2. 2.
    Identify $\mathcal{C}_{u}$ and $\mathcal{C}_{v}$ by finding their roots
  3. 3.
    If $\mathcal{C}_{u}\neq\mathcal{C}_{v}$, merge them, perform a pivoting step (cf. Section V.3) and go back to step 1.
  4. 4.
    If instead $\mathcal{C}_{u}=\mathcal{C}_{v}$, find all roots of the rigid clusters to which $u$ and $v$ belong  [^4] Note that $\mathcal{C}_{u}=\mathcal{C}_{v}$ happens only if both $u$ and $v$ are active, hence each of them belongs to at least one rigid cluster..
  5. 5.
    If a rigid cluster $\mathcal{R}$ containing both $u$ and $v$ does not already exists, i.e., if a common root has not been identified by the previous step, perform a rigidification step (cf. Section V.4) and go back to step 1.
  6. 6.
    Otherwise, perform an overconstraining step (cf. Section V.5) and go back to step 1.
The algorithm ends when all the bonds have been activated, hence covering the entire filling fraction range.
The occurrence of pivoting, rigidification or overconstraining events typically depends on the bond concentration, as we show in Fig. 2. We now describe steps 3, 5, 6 in more details.
V.3 Pivoting
At low concentration, and in particular below the CP transition, the activation of new bonds typically leads to pivoting, as shown in Fig. 2.
A pivoting step requires (i) to increase the pivotal class of both $u$ and $v$, (ii) to initialize a new $\mathcal{T}_{\rm{R}}$ tree with size one and root $uv$, (iii) to update the pebble graph by covering $uv$ with a pebble. Note that, by Theorem 1, the bond is already known to be independent, and having one pebble to cover it is therefore sufficient. Hence, if at least one node between $u$ and $v$ has a pebble, we simply use it without searching for it. If both have zero pebbles, a pebble search of type I is performed. Note that this search never fails. As the size of $\mathcal{C}_{u}$ and $\mathcal{C}_{v}$ are known, the search is performed across the smallest of the two to make it faster.
A pivoting step has constant cost if the pebble search is not performed, otherwise the cost is that of the only search performed. Overall, the cost of a pivoting step is quite low, see SM 2.
V.4 Rigidification
Above the CP transition, the probability of having a macroscopic connectivity cluster increases sharply, so that the probability of activating a bond within the same connectivity cluster is large. However, while $p<p_{RP}$, rigid clusters are small (connectivity clusters have numerous floppy modes), leading to a large probability of a rigidification event (see Fig. 2). Namely, the new independent bond will connect a pair of nodes belonging to the same connectivity cluster, but different rigid clusters, and its activation will lead to the coalescence of a variable number of distinct rigid clusters into a single one. The construction of the new rigid cluster is at the core of our algorithm.
The most obvious strategy to construct this rigid cluster is that of the JH algorithm, i.e., starting with all nodes unmarked except $u$ and $v$ that are marked rigid, use a Breadth First Search over the neighbours of nodes marked rigid. In this way, the rigid cluster is constructed by identifying a surrounding shell of nodes marked floppy. However, the implementation of this approach results in an algorithm that we observe to be quadratic (see SM 3 A).
The key of our approach is first to realize that it is sufficient to test mutual rigidity of at most one bond per existing rigid cluster, with respect to the newly activated bond. We conveniently choose these bonds as the roots of the rigid clusters. Indeed, by transitivity (Proposition 1), if the root of a rigid cluster $\mathcal{R}$ is rigid (floppy) with respect to $uv$, then all the bonds in $\mathcal{R}$ are also rigid (floppy) with respect to $uv$, and do not need to be checked. To build the rigid cluster, one could therefore test mutual rigidity between $uv$ and all the roots of the rigid clusters belonging to the connectivity cluster $\mathcal{C}$. While testing the roots in an arbitrary order does not result in satisfactory scaling (see SM 3 B), the approach becomes highly efficient when the roots are tested by following paths in a Pivot Network, as we now explain.
To construct the rigid cluster resulting from rigidification, we start a Breadth First Search from uvuv over roots of rigid clusters that could potentially coalesce into the forming one. To identify them, we introduce the Pivot Network $\mathcal{P}$, a graph whose nodes represent the rigid clusters existing in the system. If two rigid clusters share a pivot node in the physical system, they share an edge in the Pivot Network. Note that, from the “two-pivots rule” (Proposition 2), two distinct rigid clusters can share at most one pivot. In practice, the nodes of the Pivot Network correspond to the root bonds of the physical system, that uniquely identify each rigid cluster. The goal of the rigidification step is to identify the $k$ coalescing clusters/roots of Theorem 3, in order to construct $\mathcal{R}^{+}=\mathcal{R}_{1}\cup\mathcal{R}_{2}\cup\cdots\cup\mathcal{R}_{k}\cup\{uv\}$. Initially, $\mathcal{R}^{+}$ is composed of the new activated bond $uv$ only and both $u$ and $v$ are pivots of it. The BFS over the Pivot Network is performed by identifying the roots connected to $\mathcal{R}^{+}$ in the Pivot Network, and by checking the mutual rigidity of each of these roots with respect to $uv$. For each root $r$ found to be rigid with respect to $uv$, the corresponding rigid cluster is coalesced into $\mathcal{R}^{+}$, i.e., $\mathcal{R}^{+}\to\mathcal{R}^{+}\cup\mathcal{R}_{r}$. The pivots of each coalesced rigid cluster $\mathcal{R}_{r}$ become pivots of $\mathcal{R}^{+}$. The BFS then proceeds by further checking the new roots connected to $\mathcal{R}^{+}$ through these pivots. As shown in Fig. 3, this iterative procedure builds the rigid cluster resulting from rigidification by walking the paths in the Pivot Network that connect all the rigid clusters coalescing into $\mathcal{R}^{+}$. The BFS stops when a surrounding shell of floppy clusters has been identified.
Practically, the cost of the algorithm is dominated by the cost of the rigidification steps (whose cost generally depends on the bond concentration, see next section).
V.5 Overconstraining
At high bond concentration, and in particular right above the RP transition, the existence of a macroscopic rigid cluster leads to a sharp increase of the probability of overconstraining events, namely of connecting two nodes belonging to the same rigid cluster (see Fig. 2).
An overconstraining step only requires to update the $\mathcal{T}_{\rm{R}}$ of the rigid cluster to which both $u$ and $v$ belong by directing $uv$ to its root. This operation clearly has constant cost.
Figure 2: Stages of the rigidity transition. The blue, orange and green curves represent the probability, as a function of the bond concentration, that the activation of a new bond results respectively in a pivoting event, a rigidification event or an overconstraining event. We indicate the critical bond concentrations $p_{c}^{CP}=2\sin(\pi/18)$ [48] and $p_{c}^{RP}\approx 0.6602$ [32, 5] of the CP and RP transition respectively. Curves are shown for $L=2^{10}$ and are observed not to depend on $L$.
VI Implementation of the algorithm
We provide an explicit C++ implementation of the algorithm presented in the previous section for RP on the triangular lattice [49]. The generalization to any two-dimensional graph, including off-lattice settings relevant for particle systems, is straightforward.
VI.1 Data structures
The system is a triangular lattice of linear size $L$, where $i=0,\cdots,N-1$ integers are used to label nodes and $b=0,\cdots,M-1$ integers are used to label bonds, with $N=L^{2}$ and $M=3N$. $\mathcal{T}_{\rm{C}}$ and $\mathcal{T}_{\rm{R}}$ are two arrays of length $N$ and $M$ respectively. $\mathcal{T}_{\rm{C}}$ and $\mathcal{T}_{\rm{R}}$ are treated using the-non recursive implementation of path compression of Ref. [44].
The system is represented by a standard adjacency list, the pebble graph is a $2N$ array where entries $2i$ and $2i+1$ are the pebble neighbors of node $i$. Negative values correspond to inactive pebble edges. We further define several other arrays of length $N$, the use of which is detailed in SM 4, including notably an array $\pi$, which stores the pivotal class of each node. The Pivot Network $\mathcal{P}$ is implemented as an array of sets with $M$ entries. C++ sets are balanced binary search trees that guarantee that no duplicates exist and, for a set with $n$ elements, they have $\mathcal{O}(\log n)$ cost of insertion, deletion and search  [^5] We preferred sets to unordered sets as we observed the former to perform better than the latter. Indeed, the complexity of basic operations on sets is guaranteed to be $\mathcal{O}(\log n)$. In contrast, while unordered sets offer $\mathcal{O}(1)$ complexity, their performance can degrade to $\mathcal{O}(n)$ due to rehashing of the underlying hash table.. Entries corresponding to root bonds contain the pivots of the corresponding rigid cluster. All other entries are empty sets. If the activation of $uv$ triggers a pivoting step, both $u$ and $v$ are inserted in the entry of bond $uv$ if their pivotal class becomes larger than one. Since in this case $u$ and $v$ become pivots of the other rigid cluster(s) they belong to, they must be inserted in the corresponding entries of $\mathcal{P}$ as well. We identify these rigid clusters by applying find operations on the active bonds adjacent to $uv$. The rigidification step is the more involved one, and we describe it in the next section. The Pivot Network is not affected by overconstraining.
VI.2 Rigidification step
The Pivot Network is used to efficiently perform the BFS over the rigid clusters. We start the BFS with $\mathcal{R}=uv$ and iterate over its pivots, which are stored in the corresponding entry of $\mathcal{P}$. Every time two rigid clusters coalesce, $\mathcal{P}$ must be updated accordingly.
The first part of the rigidification step requires to update the pebble graph. To this aim, we (i) gather four pebbles at the new bond by performing pebble searches of type I. Note that we know the bond to be independent, so these searches never fail (implying that we never fail pebble searches of type I); (ii) Use one of the pebbles to cover the edge and update the pebble graph; (iii) Initialize a new $\mathcal{T}_{\rm{R}}$ tree with size one and root $uv$, just like in a pivoting step; (iv) Increase the pivotal class of $u$ and $v$ and update $\mathcal{P}$ accordingly. Finally, (v) freeze the remaining three pebbles at their location.
Then we proceed with the BFS, which starts by inserting $u$ and $v$ in a queue of pivots. Every time a pivot is inserted in this queue, we label it as enqueued using a dedicated array (see SM 4). All nodes are initially unmarked, except for $u$ and $v$ that are marked rigid. The rigidification step proceeds as follows:
  1. 1.
    If the queue of pivots is empty, terminate the rigidification step. Otherwise pick a pivot $p$ and remove it from the queue.
  2. 2.
    Iterate over the bonds of $p$ . For each bond, find its root $ij$. Once all the roots are processed as detailed below, go back to step 1 and process the next pivot.
  3. 3.
    Consider the mark on node $i$.
    (i) If $i$ is marked rigid →go to step 4.
    (ii) If $i$ is marked floppy →go back to step 2.
    (iii) If $i$ is unmarked, perform a pebble search of type II to mark it and then operate according to cases (i) and (ii).
  4. 4.
    Consider the mark on node $j$ and operate like in step 3.
    (i) If $j$ is marked rigid →go to step 5.
    (ii) If $j$ is marked floppy →go back to step 2.
    (iii) If $j$ is unmarked, perform a pebble search of type II to mark it and then operate according to cases (i) and (ii).
Note that the next steps are executed only if both $i$ and $j$ have been marked as rigid. This means, by transitivity, that the whole rigid cluster rooted by the bond $ij$, which we denote $\mathcal{R}_{ij}$, is mutually rigid with respect to the new bond and must be coalesced into the forming cluster $\mathcal{R}^{+}$. To this end, we first identify the largest among $\mathcal{R}^{+}$ and $\mathcal{R}_{ij}$. Let’s assume, without loss of generality, that $\mathcal{R}_{ij}\equiv\mathcal{R}_{\rm large}$ is larger than $\mathcal{R}^{+}\equiv\mathcal{R}_{\rm small}$. We perform the following steps.
  1. 5.
    Update the trees $\mathcal{T}_{\rm{R}}$: redirect the root $r_{\rm small}$ of $\mathcal{R}_{\rm small}$ to the root $r_{\rm large}$ of $\mathcal{R}_{\rm large}$ and update the entry of $r_{\rm large}$ in $\mathcal{T}_{\rm{R}}$ to keep track of the updated cluster size. This step is the analog of a standard coalescence step in the NZ algorithm for CP (see sec. III).
  2. 6.
    Remove $p$ from the set of pivots of $\mathcal{R}_{\rm small}$, stored in $\mathcal{P}_{r_{\rm small}}$ (after coalescence, $p$ is not anymore a pivot between $\mathcal{R}_{\rm large}$ and $\mathcal{R}_{\rm small}$) and decrease its pivotal class by one, $\pi_{p}\to\pi_{p}-1$. If this results in $\pi_{p}=1$ ($p$ belongs only to one cluster and is not a pivot anymore) remove $p$ from $\mathcal{P}_{r_{\rm large}}$ as well.
  3. 7.
    Iterate over the remaining pivots $p^{\prime}$ of $\mathcal{R}_{\rm small}$. If $p^{\prime}$ is already in $\mathcal{P}_{r_{\rm large}}$, $\pi_{p^{\prime}}\to\pi_{p^{\prime}}-1$. If $\pi_{p^{\prime}}$ becomes 1, remove $p^{\prime}$ from $\mathcal{P}_{r_{\rm large}}$. If $p^{\prime}$ is not yet in $\mathcal{P}_{r_{\rm large}}$, insert it and go to step 8. In any case, remove $p^{\prime}$ from $\mathcal{P}_{r_{\rm small}}$.
  4. 8.
    If $p^{\prime}$ is not labeled as enqueued, insert it in the queue of pivots and label it as enqueued. This is absolutely crucial to prevent the insertion of the same pivot in the queue more than once.
This process ends only when the queue is empty, i.e., when all the pivots have been processed. This protocol implements the rigidification strategy described in section V.4 and in Fig. 3: processing all pivots means that we attempt to build all possible paths in the Pivot Network that, starting from a rigid cluster that coalesces into $\mathcal{R}^{+}$, connect it to other existing rigid clusters. The queue is empty when the only remaining paths lead to floppy clusters.
Note that, as long as the rigidification process is not over, $\mathcal{R}_{\rm small}$ and $\mathcal{R}_{\rm large}$ may share several pivots, whose pivotal classes get updated at step 7. As can be seen in Fig. 3 (the green and red clusters in panel (c)), such pivots arise from previous coalescence events, and do not contradict Proposition 2.
Moreover, note that we only enqueue pivots of $\mathcal{R}_{\rm small}$ (step 7), namely the BFS progresses always from the pivots of the smallest cluster without enqueuing the ones of the largest cluster. This implies that, at the end of the process, the pivots of exactly one of the $\mathcal{R}_{1},\cdots\,\mathcal{R}_{k+1}$, say $\bar{\mathcal{R}}$, have not been enqueued. To see this, consider a coalescence between $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ and then a coalescence between $\mathcal{R}_{1}\cup\mathcal{R}_{2}$ and $\mathcal{R}_{3}$. After the first coalescence, the non-enqueued pivots belong to one rigid cluster only (say $\mathcal{R}_{1}$). After the second event, if $\mathcal{R}_{\rm large}=\mathcal{R}_{3}$, then the pivots of $\mathcal{R}_{1}$ gets enqueued and the ones of $\mathcal{R}_{3}$ are the only ones not enqueued, while in the opposite case the only non-enqueued ones remain those of $\mathcal{R}_{1}$. The same reasoning applies to all the successive coalescences. Ignoring the pivots of $\bar{\mathcal{R}}$ still leads to the correct construction of $\mathcal{R}^{+}$. Namely, none of the $\mathcal{R}_{1},\cdots\,\mathcal{R}_{k+1}$ is “forgotten” during rigidification. Indeed, by Theorem 3, any rigid cluster that shares a pivot with $\bar{\mathcal{R}}$ and that coalesces into $\mathcal{R}^{+}$ shares at least one other pivot with at least another coalescing cluster, from which it will be eventually reached by the BFS. In other words, clusters that share a pivot with $\bar{\mathcal{R}}$ but with none of the other coalescing ones are surely floppy. This is illustrated in Fig. 3, where the pivot of the large (red) cluster labeled $p^{*}$ is never enqueued. The green rigid cluster connected to the large one via $p^{*}$ is nonetheless reached from another pivot, and coalesced.
Not enqueuing all pivots is crucial to make the algorithm fast. Indeed, when a large rigid cluster exists, a very large number of pivots that lead to floppy clusters would be enqueued, making the iteration of step 1 exceedingly long. This is especially true above the RP transition, where the macroscopic rigid cluster has $\mathcal{O}(N)$ pivots: enqueuing them would result in a quadratic algorithm.
Figure 3: Example of a rigidification process as implemented by our algorithm. The different steps are illustrated both on the triangular lattice (top row) and on the corresponding Pivot Network (bottom row). Each color indicates a different rigid cluster. Top row: pivots are indicated with black dots. At each step, enqueued pivots are highlighted as stars. (a) The blue bond is the newly activated one: its two end nodes, marked with yellow stars, are initially enqueued in the queue of pivots. Each rigid cluster is represented by a different color. Nodes with the same color of their bonds belong to that rigid cluster only, black nodes are pivots. (b) The system state after having checked the rigid clusters led to by the pivots enqueued in panel (a). Red stars indicate pivots that have been processed. The arrows illustrate the Machta displacements $d_{p}^{r_{\rm large}}$ and $d_{p}^{r_{\rm small}}$, from pivot $p$ to the large and small roots. (c) The system state after having checked the rigid clusters led to by the pivots enqueued in panel (b). The displacement $d_{u_{s}}^{r_{\rm large}}$ has been updated according to Eq. (1). The arrows show the displacements $d_{p^{\prime}}^{r_{\rm large}}$ and $d_{p^{\prime}}^{r_{\rm large}\,^{\prime}}$ given by Eq. (2), whose difference allows to detect that wrapping has occurred along the vertical axis. Note that the pivot $p^{*}$ is never enqueued, but that the green cluster is nonetheless reached via another pivot (yellow star), cf. discussion at the end of sec. VI. (d) The system state after having checked the rigid clusters led to by the pivots enqueued in panel (c). Black stars indicate processed pivots leading to floppy clusters. (e) The system state after having checked the rigid clusters led to by the pivots enqueued in panel (d). (f) The final lattice configuration. Bottom row: the pivot network corresponding to each panel of the top row. Nodes correspond to rigid clusters, edges represent pivots. Note that pivots with pivotal class three form triangles in this network; we indicate again the non-enqueued pivot $p^{*}$. Note that as rigid clusters coalesce, nodes disappear and the network is rearranged. Also, note that the bond representing pivot $p^{*}$ disappears from panel (i) to panel (l). If it had larger pivotal class, it would still be an edge in the Pivot Network.
VII Detecting wrapping
VII.1 The Machta algorithm in CP
In systems with doubly periodic boundary conditions, the criterion for percolation is the wrapping of a cluster around one, the other, or both periodic boundaries. This corresponds to the existence of a non-contractible loop inside a cluster. A clever way to detect wrapping in CP is the Machta algorithm [42], that we recall briefly before generalizing it to the more involved RP problem (we refer to [44, 42] for a more detailed explanation of the Machta algorithm). The key difference between a contractible loop (non-wrapping) and a non-contractible one (wrapping) is the algebraic distance traveled as one goes around the loop. The essence of Machta algorithm is to exploit this fact, by keeping track of the algebraic displacements from each node to a given reference node inside its cluster, conveniently taken as the root. When a new activated bond connects two nodes of a same cluster, the difference of their respective displacements to the root is computed: wrapping has occurred if the difference of displacements is of the order of the system’s length. In the original Machta algorithm for CP, two variables are assigned to each node, that give the $x$ and $y$ displacements along a path inside the cluster, from that node to its parent in the tree. Machta displacements are updated in sync with the tree, whenever roots get redirected and when path compression is performed. Wrapping is checked only when connecting two nodes that are found to belong to the same cluster, as this is the only cluster event that might lead to wrapping.
VII.2 A Machta algorithm for RP
In RP, a first complication comes from the fact that nodes may belong to more than one rigid cluster. For each node, we therefore keep track of the displacements associated to each rigid cluster it belongs to. Namely, if node $u$ belongs to $k$ rigid clusters, so that it is shared by bonds $b_{1},\,b_{2},\cdots,b_{k}$, whose respective parents in the trees are $a_{1}=(u_{1},v_{1}),\,a_{2}=(u_{2},v_{2}),\cdots,a_{k}=(u_{k},v_{k})$, we compute the Machta displacements $d_{u}^{a_{i}}\equiv\left\{\delta_{x}(u\to u_{i}),\delta_{y}(u\to u_{i})\right\}$, for $\quad i=1,\cdots,k$. We conventionally set the reference point as the first node ($u_{i}$) of the parent bond. Two arrays of length $N$ are used to store the displacements along the two dimensions of the lattice. In our C++ implementation, each entry in the array contains a map. So, for example, to store the horizontal displacements of $u$ we use the array $dx$ such that $dx[u]$ is a map and the displacement $\delta_{x}(u\to u_{i})$ is stored as the value with key $a_{i}$, i.e, $dx[u][a_{i}]$.
The second notable difference with CP is the type of cluster events that may lead to wrapping. In RP, wrapping may occur both during overconstraining and rigidification. Below we detail the update of displacements and wrapping detection, when activating the new bond $b=(u,v)$.
Pivoting – Bond $b$ is a rigid cluster of size one (cf. sec. IV), and the root of its own tree. The displacement of its first node is therefore set as $d_{u}^{b}=\left\{0,0\right\}$ (reference node), while $d_{v}^{b}=\left\{\delta_{x}(v\to u)\right\}$.
Overconstraining – This case is akin to CP: adding a new bond to an existing rigid cluster may “close it” around the periodic boundary(ies). Therefore, when activating $b=(u,v)$ leads to overconstraining, we simply compare the total displacements to the root for $u$ and $v$.
Rigidification – Each time a rigid cluster is coalesced into the forming one (as described in section VI.2), the root of the smallest cluster, $r_{\rm small}=(u_{s},v_{s})$ is redirected to the root of the largest, $r_{\rm large}=(u_{l},v_{l})$. Denoting $p$ the pivot node between $\mathcal{R}_{\rm small}$ and $\mathcal{R}_{\rm large}$, we update the displacements as
$d_{u_{s}}^{r_{\rm large}}\to d_{u_{s}}^{r_{\rm large}\,^{\prime}}=-d_{p}^{r_{\rm small}}+d_{p}^{r_{\rm large}}$ (1)
$d_{v_{s}}^{r_{\rm large}}\to d_{v_{s}}^{r_{\rm large}\,^{\prime}}=d_{v_{s}}^{r_{\rm small}}+d_{u_{s}}^{r_{\rm large}}.$
An example update of $d_{u_{s}}^{r_{\rm large}}$ is shown in Fig. 3 (panels (b) and (c)). After coalescing $\mathcal{R}_{\rm small}$, some of its pivots are added to the pivots of $\mathcal{R}_{\rm large}$ (step 7 in sec.VI.2). We update their displacements accordingly,
$\forall p^{\prime}\in\mathcal{P}_{r_{\rm small}},\quad d_{p^{\prime}}^{r_{\rm large}\,^{\prime}}=d_{p^{\prime}}^{r_{\rm small}}+d_{u_{s}}^{r_{\rm large}}.$ (2)
To check if the coalescence step has created a non-contractible loop, we compute, for each node $p^{\prime}$ that is a pivot between small and large, the difference $d_{p^{\prime}}^{r_{\rm large}\,^{\prime}}-d_{p^{\prime}}^{r_{\rm large}}$. Namely, while in CP and overconstraining the closure of the loop is checked at the new bond, during rigidification it is more convenient to check it at pivots. This difference is the algebraic distance along the closed loop. It is zero if the rigid cluster does not (yet) wrap, non-zero if wrapping has occurred. An example of such wrapping detection is given in Fig. 3 (panel (c)). Note that the nodes of the small root may be pivots too. Therefore, we also check wrapping by computing $d_{u_{s}}^{r_{\rm large}\,^{\prime}}-d_{u_{s}}^{r_{\rm large}}$ and $d_{v_{s}}^{r_{\rm large}\,^{\prime}}-d_{v_{s}}^{r_{\rm large}}$.
VIII Results
VIII.1 Performance of the algorithm
Fig. 4 shows the performance of the algorithm presented in the previous sections, including the implementation of wrapping detection, on the triangular lattice. Quantities that depend on the filling fraction are convoluted as in [40] and shown as a function of $p$. Averages are made over $10^{5}$ for the smallest sizes, and over about $600$ for the largest one. The performance of the algorithm, quantified by the average time to complete a simulation, is observed to scale as $N^{1.02}$ (panel (a)). A striking feature of the algorithm is the extremely low number of pebble searches needed to update the system’s state after the activation of a bond, shown as a function of the bond concentration $p$ (panel (b)). This function peaks near $p_{c}^{RP}$ with a maximum that is of the order of 10 even at large system sizes. The average time of each bond activation as a function of the bond concentration is shown on panel (c). Despite this function also shows a peak near the rigidity transition, the fundamental observation here is that the time per bond is not proportional to the number of pebble searches. Therefore, contrary to rigidification protocols based on the JH approach (cf. SM 3 A), the total execution time cannot be deduced from the number of searches. For example, for $p_{c}^{CP}\lesssim p<p_{c}^{RP}$ the number of searches grows much more rapidly than the time per bond, which remains low. A deeper investigation is required to understand the scaling of the algorithm.
Figure 4: Performance of the algorithm. (a) The average time (in minutes) to complete a simulation as a function of the system size. The blue solid line is the best fit to the log-log of the data and shows the scaling $N^{1.02}$. Error bars are smaller than the markers’ size. (b) Average number of pebble searches (of any type) needed at each bond activation, as function of the bond concentration $p$, for different system sizes $N=L^{2}$. (c) Average time (in seconds) of each bond activation as function of $p$, for the same system sizes as in panel (b).
To this aim, we first note that Fig. 4a shows the performance of our code, which is only a proxy of the computational complexity of the algorithm and, in particular, sets an upper bound on it (see SM 5). The computational complexity is truly quantified by the number of operations performed. To estimate it, we focus on the rigidification step, which is the main contributor to the cost of the algorithm. The observation is that there are mainly three operations that determine how expensive a rigidification step is. The first operation is the preliminary search of four pebbles using type I searches. The second operation is the repeated application of type II searches to mark nodes. The last expensive operation is the iteration of pivots $p^{\prime}$, at step 7 of the algorithm. Note that the number of pivots $p$ in step 1 is at most equal to the number of pivots $p^{\prime}$, as only a fraction of the $p^{\prime}$ gets enqueued and hence iterated over in step 1 (with the exception of $u$ and $v$, which represent however an $N$-independent cost). To quantify how expensive pebble searches are, we count the total number of nodes visited during each search and we further count the number of pivots $p^{\prime}$. Results are shown in Fig. 5.
Figure 5: Average number of nodes visited during type I searches (blue circles, scaling as $N^{1.02}$), average number of nodes visited during type II searches (red squares, scaling as $N^{1.01}$) and average number of pivots $p^{\prime}$ over wich we iterate during step 7 of the algorithm (green triangles, scaling as $N^{1.00}$). Data are shown as a function of the system size.
Data show that the main cost of the algorithm is due to pebble searches of type I, despite they are less numerous than pebble searches of type II (which, however, are only slightly less expensive than the former) The number of pivots $p^{\prime}$ represents instead a minor contribution to the cost of the algorithm. Importantly, pebbles searches of type I are estimated to have a cost scaling as $N^{1.02}$, i.e., the same scaling inferred for the performance. This suggests that the observed performance well quantifies the computational complexity and that its (extremely small) superlinearity is an intrinsic feature of the algorithm, i.e., that it is implementation independent. Overall, we conclude that the computational complexity of the algorithm presented in this work is $N^{1.02}$, i.e., the phase diagram of the RP transition can be computed in a time that –to any practical purpose– is linear in the system size (see SM 5).
VIII.2 Estimation of the RP critical exponents and critical threshold
We apply our algorithm to the determination of the main critical exponents – the correlation length exponent $\nu$, the fractal dimension $D_{f}$, the susceptibility exponent $\gamma$ and the order parameter exponent $\beta$. We also determine the critical point $p_{c}$. We perform simulations for system sizes ranging from $L=2^{6}=64$ to $L=2^{14.5}=23170$, and choose the numbers of samples so as to activate at least $7\cdot 10^{11}$ bonds per value of $L$ (see SM 6). For each run of the algorithm, and for both the connectivity and rigid clusters, we store: the number of bonds $m_{x}$ and $m_{xy}$ at which wrapping happens for the first time, respectively along the horizontal and along both horizontal and vertical directions; the size $S_{xy}$ of the $xy$-wrapping cluster right at wrapping; the maximum value $\chi_{\rm max}$ of the susceptibility.
We then perform scaling analysis to determine $D_{f}$, $\nu$, $\gamma/\nu$ and $p_{c}$, for both the RP and CP transitions, the latter being used to benchmark our scaling analysis. The values of all exponents are reported in Table 1, giving a comparison of the CP and RP universality classes. These values are validated by the excellent collapse of the rescaled scaling functions (order parameter, susceptibility and wrapping probability) shown in Fig. 7. Overall, our results strengthen the long-standing hypothesis [5] that 2d CP and central-force RP belong to different universality classes.
VIII.2.1 Fractal dimension $D_{f}$
The fractal dimension $D_{f}$ is estimated from the (rescaled) size of the wrapping cluster $S_{xy}/(3L^{2})$, expected to scale as $S_{xy}\sim L^{D_{f}-2}$ [1]. For CP, a simple power-law fit gives an estimate in excellent agreement with the exact value $91/48$, reported in Figure 6 (a). For RP, deviations to scaling are present at small sizes, so that we choose to fit $S_{xy}^{\rm RP}$ as a power-law, but only for sufficiently large $L$. The number of datapoints removed is chosen as the one maximizing the goodness of fit, and indicated in Fig. 6. Note that we also tried to fit the data of $S_{xy}$ (and similarly for $\Delta$, $\chi_{\rm max}$ and $p_{c}$) to the form $\mathcal{O}(L)=A\,L^{y}\left(1+B/L^{w}\right)$, which incorporates corrections to scaling. For all observables, such fits lead values of the exponent $y$ consistent with those obtained by removing the smallest sizes from a power-law fit.
VIII.2.2 Susceptibility exponent $\gamma$
The maximum of the susceptibility is known to scale as $\chi_{max}\sim L^{\gamma/\nu}$ [1], from which we estimate the exponent $\gamma/\nu$. As for the size of the wrapping cluster $S_{xy}$ (cf. section VIII.2.1), the data for CP is perfectly fitted by a power-law, while RP shows significant deviations to scaling at small sizes, see Figure 6 (b). We adopt the same strategy as in section VIII.2.1 and fit only data with sufficiently large $L>L_{\min}$, the value of $L_{\rm min}$ being chosen as maximizing the goodness of fit. The susceptibility exponent $\gamma^{\rm RP}$ is then computed using the value of $\nu^{\rm RP}$ obtained independently in section VIII.2.3, and reported in Table 1.
VIII.2.3 Correlation length exponent $\nu$
The measurements of $m_{x}$ and $m_{xy}$ allows to compute the wrapping probabilities $R_{x}$ (along the horizontal axis), $R_{xy}$ (along both axes), $R_{e}$ (along either axis) and $R_{1}$ (along one axis), see [44] and Fig. S6. For each wrapping probability, we compute the width of the transition region $\Delta(L)$, as explained in SM 6 A. For each wrapping type, $\Delta(L)$ is expected to scale as $\Delta(L)\sim L^{-1/\nu}$ [1].
For CP, simple power-law fits of $\Delta(L)$ yield exponents in excellent agreement with the exact value $\nu^{\rm CP}=4/3$, for all four wrapping types. For RP, small deviations to scaling are visible in the residual plots, especially for $R_{x}$ and $R_{xy}$, while they appear suppressed in $R_{e}$ and $R_{1}$. We adopt the same strategy as in section VIII.2.1 and fit the data for $\Delta(L)$ as power-laws, discarding the smallest sizes as indicated in Figure 6 (c). The corresponding values of $1/\nu^{\rm RP}$ are reported in Figure 6 (c). Note that not all four wrapping probabilities are independent [40]. We find that the estimates from $R_{x}$ and $R_{xy}$ lead to the best collapse for all four wrapping probabilities, and therefore compute the final value of $\nu^{\rm RP}$, reported in Table 1, as the average of $\nu^{\rm RP}_{x}$ and $\nu^{\rm RP}_{xy}$.
Figure 6: Scaling analysis for RP. (a), (b) Size $S_{xy}/(3L^{2})$ of the wrapping cluster, and maximum $\chi_{\rm max}$ of the susceptibility, both rescaled by the total number of bonds, as function of $L$, shown for CP and RP. The lines are power-law fits of the data marked with filled symbols. (c) Widths $\Delta(L)$ of the transition region for each wrapping type x,xy,e, 1x,\,xy,\,e,\,1, as functions of the system size $L$. The slopes of the solid lines are reported in the legend and are obtained from power-law fits of the data marked with filled symbols. For the sake of readability, data from $R_{xy}$, $R_{e}$ and $R_{1}$ have been shifted by a factor 2, 4 and 6 respectively. (d) Finite-size critical thresholds $p_{c}/\Delta$, for each wrapping type, as functions of $\Delta^{-1}$. The slopes of the solid lines, reported in the legend, give the corresponding values of $p_{c}^{\rm RP}$. Data has been shifted and shown in log-log scale to ease readability.
VIII.2.4 Critical threshold $p_{c}$
The wrapping probabilities $R_{i}$ allow to further compute, for each wrapping type, the size-dependent transition thresholds $p_{c}(L)$, namely the bond concentration at which the emergence of a wrapping cluster is expected on a finite lattice of size $L$ (cf. SM 6 A). In general we expect $p_{c}(L)-p_{c}\propto L^{-1/\nu}\propto\Delta(L)$, where $p_{c}=p_{c}(L\to\infty)$. For RP, pronounced deviations to scaling are visible at small sizes, namely $p_{c}^{\rm RP}(\Delta)=p_{c}^{\rm RP}+A\,\Delta^{\rm RP}\left(1+B\,\Delta^{w}\right)$ for some positive exponent $w$. We consider the quantity $p_{c}^{\rm RP}/\Delta^{\rm RP}$ and fit
$y(x)\equiv\frac{p_{c}^{\rm RP}(\Delta^{\rm RP})}{\Delta^{\rm RP}}=p_{c}\;x+A$ (3)
for $L>L_{\rm min}$ sufficiently large, namely when $x=1/\Delta^{\rm RP}$ is sufficiently large that the correction term $\propto x^{-w}$ becomes negligible. For each wrapping type the value of $L_{\rm min}$ is chosen to maximize the goodness of fit. All four wrapping types yield excellent fits, as shown in Fig. 6 (d). Our final estimate of $p_{c}^{\rm RP}$ reported in Table 1 is the average of the estimates obtained from $R_{x}$ and $R_{xy}$, that have the lowest uncertainties.
Figure 7: Scaling functions for RP (main plots), and their collapse (insets) when rescaled using the values of the critical threshold and exponents in Table 1. (a) Order parameter $P_{\infty}$ (b) Susceptibility $\chi$ (c) Wrapping probability $R_{xy}$.
CP RP
$D_{f}$ $91/48\approx 1.895833$ $1.8423(7)$
$\nu$ $4/3\approx 1.333333$ $1.1694(8)$
$\gamma$ $43/18\approx 2.388889$ $1.928(3)$
$\beta$ $5/36\approx 0.138889$ $0.1844(8)$
$p_{c}$ $2\sin(\pi/18)\approx 0.347296$ $0.6602741(4)$
Table 1: CP and RP universality classes, and critical thresholds. The value of $\beta$ is computed using hyperscaling as $\beta=\nu(2-D_{f})$.
IX Conclusion
In this work we have presented novel theoretical results on two dimensional central-force Rigidity Percolation and have exploited them to derive an exact algorithm that computes the whole phase diagram of the model in a time that is almost linear in the system size. A C++ implementation of the algorithm is provided in [49]. We have assessed the complexity intrinsic to the algorithm on the triangular lattice, where each node has exactly six neighbors. The rigidity transition, however, is of interest also on different topologies [26, 27]. Our algorithm can be straightforwardly implemented on any system where rigidity is ruled by Laman theorem. In particular, it can be readily used to analyze the rigidity of off-lattice configurations obtained from simulations of particle systems. On networks with a broad degree distribution, where the average degree might grow significantly with $N$ [51], the repeated application of the find operation would introduce a supplementary, $N$-dependent cost. In any other network as, e.g., Erdős–Rényi graphs we expect our algorithm to perform well. The actual performance might depend on fine details of the system such as the structure of the pebble graph (pebble searches might be more or less expensive than on the triangular lattice) and on the geometry of rigid clusters (number of pivots they have).
Our algorithm pushes the size limitation encountered in previous works [5, 6, 30, 31], allowing us to characterize accurately the universality class: we have performed scaling analysis on systems up to 536 848 900536\,848\,900 nodes, and provide new estimates of the main critical exponents, $\nu$, $\beta$, $D_{f}$ and $\gamma$, as well as the critical threshold $p_{c}$. The consistency of these values is validated by the excellent collapse of the rescaled scaling functions, shown in Fig. 7.
On the theoretical side, our results provide new material to the long-standing question of the relationship between CP and RP [52, 23, 16], by showing explicitly that different cluster merging mechanisms are at play in the two problems. RP features notably non-local “rigidification” events that lead to the coalescence of several (rigid) clusters, a phenomenon absent in CP. In contrast with the standard Pebble Game, we construct the system in a dynamic way, making it straightforward to study accurately the frequency and spatial extension of rigidification events  [^6] While concluding the first version of this work we became aware of the preprint [61], that proposes a similar dynamic approach to RP. The algorithm has not been published, but is reported to perform slightly worse than the standard Pebble Game.. It would also be very interesting to generalize this approach to study rigidity loss [54], in particular large-scale fluidization events whereby the removal of a single bond makes a macroscopic rigid cluster collapse into smaller clusters connected by non-rigid hinges.
Although we focused here on networks with central forces, it is important to examine if some of our theoretical conclusions generalize to systems with non-central interactions. This could allow to combine generalized pebbles games [15, 33] with the NZ approach we presented to develop faster algorithm for RP with, for instance, frictional interactions. Furthermore, our exact results may be exploited to inform non-exact, neural network-based algorithms for rigidity transitions [55]. Finally, the exact determination of rigid clusters in three (or more) dimensions remains a crucial bottleneck [22, 20]. Very recent results [56] have opened the possibility of designing exact algorithms for rigidity analysis in any dimension, and we can hope to make these highly efficient by incorporating the insights of the present work.
Note added to the second version: While finalizing the second version of this work –which completes our original study by determining the critical exponents, we became aware of the preprint [57], which also reports new estimates for the RP critical exponents. To the best of our knowledge (the code not yet being publicly available), the algorithm described in [57] incorporates several strategies that we introduced in the first version of our work, but the system sizes considered in their analysis do not exceed $L=8192$. We observe that the reported values of the critical threshold and the correlation length exponent are in good agreement with our results. The larger value of the fractal dimension, $D_{f}=1.850(2)$, may be explained by the fact that it is obtained from a different set of observables, wrapping not being implemented in their algorithm.
X Acknowledgments
This research was funded in whole by the Austrian Science Fund (FWF) under Project. No. Y-1163-N27. For open access purposes, the author has applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. DN acknowledges computation time at the Vienna Scientific Cluster (VSC). NJ has been supported by a grant from MIAI@Grenoble Alpes (ANR-19-P3IA-0003). Part of the computations presented in this paper were performed using the GRICAD infrastructure (https://gricad.univ-grenoble-alpes.fr), which is supported by Grenoble research communities.
The authors are grateful to Lorenzo Rovigatti, Silvano Ferrari, Emanuela Bianchi and Claudio Castellano for insightful discussions and valuable comments on the manuscript. We especially thank Hugo Vanneuville and Vincent Beffara for their precious help in finalizing the proofs given in the SM. Finally, we thank the organizers of the 2024 Network Dynamics workshop in Les Houches, where this project started.
XI Code availability
The code is available at [49].